To finance some manufacturing tools it needs for the next 3 years, Waldrop Corporation is considering a leasing arrangement. The tools will be obsolete and worthless after 3 years. The firm will depreciate the cost of the tools on a straight-line basis over their 3-year life. It can borrow $4,900,000, the purchase price, at 8% and buy the tools, or it can make 3 equal end-of-year lease payments of $2,050,000 each and lease them. The loan obtained from the bank is a 3-year simple interest loan, with interest paid at the end of the year. The firm's tax rate is 30%. Annual maintenance costs associated with ownership are estimated at $220,000, but this cost would be borne by the lessor if it leases. What is the net advantage to leasing (NAL), in thousands? (Suggestion: Delete 3 zeros from dollars and work in thousands.)

$96

$106

$118

$130

$138
Answer #1

OPTION 1: PURCHASE OPTION

The loan amount is repayable with interest @ 8% in 3 equal annual installments at the end of each year.
The PVAF @8% for 3 years is 2.577
Therefore Annual Installment =  $4,900,000 / 2.577 = $1,901,364

We have to calculate the Interest cost for each year, for which we will first make a schedule of Debt Repayment which is attached below:
Schedule of Debt Repayment (in thousand $) End of Total Payment Principal Year (Installment) Interest Principal Outstanding 4
Note: Year 3 Interest is the Balancing figure.

Now we will calculate the Present Value of Total Outflows from Purchase (Debt) Alternative, for which we will make a schedule of Cash Outflows as attached below:
End of Year Annual Installment (a) 1,901.36 1,901.36 1,901.36 Schedule of Cash Outflows: Purchase/Debt Alternative in thousan

Therefore, Total Present Value of Outflows under Debt/Purchase Option = $ 3,995.29 thousand

Note: While Discounting the Total Cash Outflow, we will use the after tax discounting rate, which is
8 * (1 - 0.3) = 8 * 0.7 = 5.6%


OPTION 2: LEASE OPTION

Lease Rent = $ 2,050 thousand
Tax Shield on above = 2,050,000 * 0.3 = $ 615 thousand

Therefore, Annual Cash Outflow = 2,050 - 615 = $ 1,435 thousand
After Tax Cost of debt = 8 * (1 - 0.3) = 8 * 0.7 = 5.6%
PVAF @ 5.6% for 3 years = 2.693

Therefore, PV = $ 1,435 * 2.693 = $  3,864.34 thousand.
Therefore, Total Present Value of Outflows under Lease Option = $ 3,864.34 thousand

Therefore,
NAL = $ 3,995.29 thousand - $ 3,864.34 thousand
= $ 130.95 thousands


Therefore, correct alternative from the given options is $ 130

Note : Ignore 0.95 as rounding off error